# TabImpute: Accurate and Fast Zero-Shot Missing-Data Imputation with a Pre-Trained Transformer

### v1

@@ A.3 Additional tables
Table 4: Imputation Accuracy ± Standard Deviation by Missingness Pattern (Remaining Methods)
Pattern K-Nearest Neighbors ICE Col Mean TabPFN
MCAR 0.762 ± 0.158 0.658 ± 0.273 0.473 ± 0.263 0.410 ± 0.257
Col-MAR 0.846 ± 0.183 0.832 ± 0.248 0.627 ± 0.299 0.553 ± 0.330
NN-MNAR 0.757 ± 0.168 0.643 ± 0.254 0.507 ± 0.266 0.437 ± 0.266
Block-MNAR 0.857 ± 0.184 0.813 ± 0.239 0.774 ± 0.241 0.572 ± 0.335
Seq-MNAR 0.889 ± 0.114 0.819 ± 0.240 0.778 ± 0.214 0.641 ± 0.290
Self-Masking-MNAR 0.623 ± 0.268 0.681 ± 0.328 0.292 ± 0.258 0.268 ± 0.248
Panel-MNAR 0.614 ± 0.400 0.584 ± 0.407 0.502 ± 0.389 0.442 ± 0.384
Polarization-MNAR 0.631 ± 0.199 0.547 ± 0.279 0.972 ± 0.059 0.914 ± 0.187
Soft-Polarization-MNAR 0.575 ± 0.266 0.723 ± 0.217 0.754 ± 0.207 0.598 ± 0.327
Latent-Factor-MNAR 0.776 ± 0.186 0.716 ± 0.268 0.601 ± 0.297 0.508 ± 0.316
Cluster-MNAR 0.784 ± 0.150 0.758 ± 0.279 0.559 ± 0.278 0.500 ± 0.297
Two-Phase-MNAR 0.904 ± 0.124 0.881 ± 0.208 0.666 ± 0.273 0.466 ± 0.329
Overall 0.751 ± 0.238 0.721 ± 0.289 0.625 ± 0.313 0.526 ± 0.334
Pattern MIWAE MICE GAIN
MCAR 0.236 ± 0.274 0.316 ± 0.283 0.543 ± 0.256
Col-MAR 0.491 ± 0.315 0.521 ± 0.354 0.276 ± 0.354
NN-MNAR 0.234 ± 0.280 0.310 ± 0.325 0.400 ± 0.324
Block-MNAR 0.594 ± 0.267 0.491 ± 0.311 0.184 ± 0.317
Seq-MNAR 0.586 ± 0.289 0.478 ± 0.329 0.204 ± 0.311
Self-Masking-MNAR 0.209 ± 0.269 0.631 ± 0.348 0.456 ± 0.380
Panel-MNAR 0.374 ± 0.363 0.251 ± 0.325 0.279 ± 0.360
Polarization-MNAR 0.535 ± 0.193 0.143 ± 0.230 0.264 ± 0.283
Soft-Polarization-MNAR 0.546 ± 0.239 0.187 ± 0.245 0.432 ± 0.409
Latent-Factor-MNAR 0.358 ± 0.296 0.331 ± 0.312 0.301 ± 0.305
Cluster-MNAR 0.321 ± 0.277 0.317 ± 0.287 0.356 ± 0.311
Two-Phase-MNAR 0.521 ± 0.321 0.557 ± 0.313 0.218 ± 0.320
Overall 0.417 ± 0.314 0.378 ± 0.338 0.326 ± 0.343
Table 5: Imputation Accuracy ± Standard Deviation by Method (Zero-shot models)
Note that the normalized numbers are slightly different from Tab. 1 because we did not include every other zero-shot method in the normalization in Tab. 1.
Method MCAR NN-MNAR Self-Masking-MNAR Col-MAR
TabImpute+ 0.749 ± 0.246 0.847 ± 0.200 0.929 ± 0.178 0.883 ± 0.255
EWF-TabPFN 0.746 ± 0.247 0.841 ± 0.201 0.935 ± 0.179 0.882 ± 0.254
TabImpute 0.860 ± 0.222 0.828 ± 0.230 0.246 ± 0.254 0.551 ± 0.350
TabImpute (MCAR then MAR) 0.751 ± 0.284 0.666 ± 0.355 0.208 ± 0.243 0.499 ± 0.354
TabImpute (More Heads) 0.690 ± 0.261 0.700 ± 0.233 0.227 ± 0.255 0.523 ± 0.321
EWF-TabPFN Fine-Tuned 0.306 ± 0.189 0.279 ± 0.220 0.406 ± 0.313 0.339 ± 0.338
TabPFN 0.066 ± 0.193 0.067 ± 0.194 0.162 ± 0.224 0.160 ± 0.254
Method Block-MNAR Seq-MNAR Overall
TabImpute+ 0.804 ± 0.272 0.672 ± 0.361 0.814 ± 0.270
EWF-TabPFN 0.795 ± 0.279 0.670 ± 0.358 0.811 ± 0.271
TabImpute 0.778 ± 0.288 0.834 ± 0.214 0.683 ± 0.342
TabImpute (MCAR then MAR) 0.722 ± 0.287 0.805 ± 0.277 0.608 ± 0.362
TabImpute (More Heads) 0.668 ± 0.258 0.729 ± 0.260 0.589 ± 0.317
EWF-TabPFN Fine-Tuned 0.385 ± 0.305 0.347 ± 0.326 0.344 ± 0.288
TabPFN 0.048 ± 0.138 0.126 ± 0.262 0.105 ± 0.218
@@
Dataset TabImpute+ HyperImpute MissForest OT
EgyptianSkulls 0.917 0.961 0.946 0.966
FacultySalaries 0.693 0.811 0.567 0.572
Glass-Classification 0.937 0.810 0.886 0.902
ICU 1.045 1.137 1.112 1.045
MercuryinBass 1.373 1.377 1.343 1.359
Rainfall-in-Kerala-1901-2017 0.942 0.902 0.950 0.958
SMSA 0.813 0.939 0.786 0.807
SolarPower 0.801 0.873 0.888 0.914
Student-Scores 0.391 0.441 0.435 0.477
Swiss-banknote-conterfeit-detection 0.976 0.856 0.730 0.786
WineDataset 0.928 0.803 0.797 0.789
alcohol-qcm-sensor 0.590 0.580 0.517 0.560
analcatdata_election2000 0.544 0.662 0.615 0.722
analcatdata_gviolence 0.795 0.801 0.877 0.794
analcatdata_olympic2000 1.867 1.922 1.895 1.945
appendicitis_test 0.860 0.650 0.792 0.753
appendicitis_test_edsa 0.668 0.532 0.569 0.586
autoPrice 0.834 0.762 0.702 0.734
baskball 1.027 1.100 1.109 1.017
benzo32 1.055 1.055 1.053 0.950
breast-cancer-coimbra 1.058 1.183 1.098 1.104
chatfield_4 0.475 0.531 0.492 0.522
chscase_vine1 0.808 0.996 0.833 0.945
dataset_analcatdata_creditscore 1.063 1.131 1.122 1.065
divorce_prediction 0.525 0.528 0.467 0.493
edm 0.763 0.553 0.566 0.570
humans_numeric 0.948 0.970 0.991 1.011
machine_cpu 1.037 0.956 0.858 0.962
metafeatures 2.263 2.090 2.225 2.263
pollution 1.967 2.065 1.837 1.945
pwLinear 1.075 1.356 1.356 1.092
pyrim 0.877 0.863 0.859 0.842
rabe_131 0.790 1.003 1.031 0.997
rabe_148 0.984 1.017 0.958 1.045
rabe_265 1.131 1.298 1.090 1.106
sleuth_case1201 1.054 0.939 0.963 0.998
sleuth_ex1605 1.087 1.257 1.071 0.969
slump 0.871 0.650 0.805 0.787
treepipit 1.264 1.169 1.172 1.167
visualizing_hamster 0.872 0.905 0.750 0.774
wisconsin 0.777 0.685 0.783 0.817
witmer_census_1980 0.704 0.616 0.637 0.653

### v3

@@ A.3 Additional tables
Table 6: Imputation Accuracy ± Standard Error by Missingness Pattern (All Methods)
Pattern TabImpute EWF-TabPFN HyperImpute OT MissForest
MCAR 0.882 ± 0.021 0.817 ± 0.024 0.804 ± 0.031 0.855 ± 0.019 0.867 ± 0.022
NN-MNAR 0.900 ± 0.019 0.876 ± 0.020 0.760 ± 0.036 0.814 ± 0.021 0.818 ± 0.029
Self-Masking-MNAR 0.703 ± 0.038 0.682 ± 0.042 0.689 ± 0.039 0.568 ± 0.037 0.638 ± 0.040
Col-MAR 0.877 ± 0.030 0.855 ± 0.027 0.814 ± 0.040 0.750 ± 0.040 0.773 ± 0.039
Block-MNAR 0.890 ± 0.026 0.905 ± 0.026 0.873 ± 0.027 0.857 ± 0.024 0.860 ± 0.023
Seq-MNAR 0.920 ± 0.014 0.905 ± 0.015 0.864 ± 0.030 0.894 ± 0.014 0.829 ± 0.031
Panel-MNAR 0.915 ± 0.025 0.711 ± 0.048 0.865 ± 0.033 0.887 ± 0.025 0.912 ± 0.020
Polarization-MNAR 0.804 ± 0.024 0.879 ± 0.021 0.622 ± 0.039 0.802 ± 0.020 0.567 ± 0.033
Soft-Polarization-MNAR 0.759 ± 0.045 0.864 ± 0.024 0.711 ± 0.034 0.750 ± 0.029 0.667 ± 0.041
Latent-Factor-MNAR 0.891 ± 0.021 0.883 ± 0.018 0.775 ± 0.036 0.842 ± 0.026 0.834 ± 0.023
Cluster-MNAR 0.903 ± 0.014 0.871 ± 0.021 0.839 ± 0.028 0.828 ± 0.022 0.830 ± 0.021
Two-Phase-MNAR 0.886 ± 0.026 0.851 ± 0.029 0.865 ± 0.031 0.807 ± 0.028 0.873 ± 0.020
Censoring-MNAR 0.710 ± 0.032 0.593 ± 0.045 0.786 ± 0.041 0.589 ± 0.036 0.671 ± 0.038
Overall 0.849 ± 0.008 0.823 ± 0.009 0.790 ± 0.010 0.788 ± 0.009 0.780 ± 0.009
Pattern K-Nearest Neighbors ICE SoftImpute Col Mean DiffPuter
MCAR 0.750 ± 0.025 0.659 ± 0.041 0.617 ± 0.044 0.478 ± 0.041 0.478 ± 0.041
NN-MNAR 0.741 ± 0.026 0.627 ± 0.039 0.679 ± 0.035 0.497 ± 0.041 0.497 ± 0.041
Self-Masking-MNAR 0.614 ± 0.038 0.673 ± 0.049 0.396 ± 0.044 0.293 ± 0.038 0.293 ± 0.038
Col-MAR 0.839 ± 0.028 0.825 ± 0.038 0.703 ± 0.040 0.622 ± 0.046 0.598 ± 0.047
Block-MNAR 0.860 ± 0.028 0.812 ± 0.037 0.799 ± 0.042 0.782 ± 0.034 0.806 ± 0.029
Seq-MNAR 0.885 ± 0.018 0.815 ± 0.037 0.783 ± 0.036 0.775 ± 0.033 0.775 ± 0.033
Panel-MNAR 0.899 ± 0.028 0.797 ± 0.044 0.340 ± 0.053 0.787 ± 0.037 0.811 ± 0.032
Polarization-MNAR 0.641 ± 0.028 0.560 ± 0.041 0.665 ± 0.048 0.964 ± 0.012 0.964 ± 0.012
Soft-Polarization-MNAR 0.580 ± 0.042 0.729 ± 0.031 0.677 ± 0.033 0.747 ± 0.030 0.747 ± 0.030
Latent-Factor-MNAR 0.764 ± 0.029 0.703 ± 0.042 0.714 ± 0.039 0.590 ± 0.045 0.590 ± 0.045
Cluster-MNAR 0.774 ± 0.024 0.750 ± 0.042 0.627 ± 0.046 0.558 ± 0.044 0.558 ± 0.044
Two-Phase-MNAR 0.898 ± 0.020 0.876 ± 0.032 0.678 ± 0.047 0.671 ± 0.042 0.671 ± 0.042
Censoring-MNAR 0.515 ± 0.047 0.866 ± 0.035 0.455 ± 0.054 0.346 ± 0.040 0.307 ± 0.037
Overall 0.751 ± 0.010 0.746 ± 0.011 0.626 ± 0.013 0.624 ± 0.013 0.623 ± 0.013
Pattern TabPFN MICE MIWAE GAIN
MCAR 0.415 ± 0.041 0.329 ± 0.044 0.254 ± 0.043 0.545 ± 0.039
NN-MNAR 0.428 ± 0.041 0.300 ± 0.049 0.230 ± 0.043 0.392 ± 0.049
Self-Masking-MNAR 0.268 ± 0.037 0.611 ± 0.051 0.217 ± 0.041 0.452 ± 0.056
Col-MAR 0.549 ± 0.050 0.520 ± 0.054 0.488 ± 0.048 0.275 ± 0.054
Block-MNAR 0.582 ± 0.051 0.502 ± 0.048 0.606 ± 0.039 0.202 ± 0.051
Seq-MNAR 0.638 ± 0.045 0.475 ± 0.050 0.584 ± 0.045 0.204 ± 0.048
Panel-MNAR 0.234 ± 0.042 0.553 ± 0.048 0.269 ± 0.045 0.536 ± 0.059
Polarization-MNAR 0.907 ± 0.029 0.192 ± 0.037 0.562 ± 0.027 0.294 ± 0.045
Soft-Polarization-MNAR 0.600 ± 0.050 0.219 ± 0.038 0.548 ± 0.036 0.425 ± 0.062
Latent-Factor-MNAR 0.499 ± 0.048 0.338 ± 0.047 0.355 ± 0.045 0.296 ± 0.046
Cluster-MNAR 0.499 ± 0.047 0.327 ± 0.044 0.330 ± 0.044 0.357 ± 0.048
Two-Phase-MNAR 0.478 ± 0.049 0.564 ± 0.048 0.529 ± 0.049 0.228 ± 0.049
Censoring-MNAR 0.342 ± 0.041 0.648 ± 0.050 0.308 ± 0.045 0.320 ± 0.051
Overall 0.495 ± 0.014 0.429 ± 0.014 0.407 ± 0.013 0.348 ± 0.015
Table 7: Mean Column-wise 1-Wasserstein Distance ± Standard Error by Missingness Pattern
Pattern ICE HyperImpute MissForest K-Nearest Neighbors MICE
MCAR 0.940 ± 0.013 0.919 ± 0.019 0.825 ± 0.019 0.742 ± 0.017 0.814 ± 0.033
NN-MNAR 0.918 ± 0.019 0.925 ± 0.020 0.852 ± 0.020 0.723 ± 0.019 0.819 ± 0.030
Self-Masking-MNAR 0.788 ± 0.035 0.719 ± 0.037 0.641 ± 0.036 0.606 ± 0.034 0.727 ± 0.054
Seq-MNAR 0.827 ± 0.025 0.842 ± 0.020 0.814 ± 0.019 0.785 ± 0.021 0.853 ± 0.030
Polarization-MNAR 0.616 ± 0.039 0.648 ± 0.042 0.540 ± 0.039 0.718 ± 0.032 0.300 ± 0.041
Soft-Polarization-MNAR 0.787 ± 0.038 0.757 ± 0.039 0.750 ± 0.045 0.529 ± 0.046 0.434 ± 0.043
Latent-Factor-MNAR 0.930 ± 0.016 0.912 ± 0.023 0.862 ± 0.019 0.735 ± 0.018 0.821 ± 0.032
Cluster-MNAR 0.919 ± 0.018 0.913 ± 0.021 0.830 ± 0.020 0.736 ± 0.017 0.811 ± 0.037
Two-Phase-MNAR 0.831 ± 0.030 0.795 ± 0.030 0.758 ± 0.023 0.835 ± 0.021 0.771 ± 0.041
Overall 0.840 ± 0.010 0.826 ± 0.011 0.764 ± 0.011 0.712 ± 0.010 0.706 ± 0.016
Pattern ForestDiffusion OT TabImpute EWF-TabPFN SoftImpute
MCAR 0.822 ± 0.021 0.721 ± 0.017 0.724 ± 0.029 0.520 ± 0.035 0.465 ± 0.060
NN-MNAR 0.756 ± 0.017 0.697 ± 0.018 0.695 ± 0.031 0.541 ± 0.034 0.400 ± 0.064
Self-Masking-MNAR 0.673 ± 0.035 0.528 ± 0.029 0.640 ± 0.035 0.601 ± 0.039 0.364 ± 0.048
Seq-MNAR 0.905 ± 0.018 0.739 ± 0.019 0.621 ± 0.029 0.576 ± 0.038 0.624 ± 0.048
Polarization-MNAR 0.267 ± 0.045 0.865 ± 0.024 0.866 ± 0.023 0.862 ± 0.027 0.640 ± 0.041
Soft-Polarization-MNAR 0.507 ± 0.036 0.662 ± 0.036 0.632 ± 0.040 0.747 ± 0.034 0.561 ± 0.042
Latent-Factor-MNAR 0.747 ± 0.022 0.707 ± 0.019 0.676 ± 0.032 0.529 ± 0.040 0.459 ± 0.059
Cluster-MNAR 0.757 ± 0.024 0.704 ± 0.018 0.663 ± 0.031 0.548 ± 0.040 0.397 ± 0.063
Two-Phase-MNAR 0.909 ± 0.025 0.678 ± 0.029 0.742 ± 0.029 0.643 ± 0.040 0.597 ± 0.046
Overall 0.705 ± 0.014 0.700 ± 0.009 0.695 ± 0.011 0.619 ± 0.013 0.501 ± 0.018
Pattern MIWAE GAIN Col Mean TabPFN
MCAR 0.420 ± 0.018 0.600 ± 0.034 0.024 ± 0.008 0.067 ± 0.016
NN-MNAR 0.414 ± 0.014 0.491 ± 0.036 0.031 ± 0.015 0.040 ± 0.013
Self-Masking-MNAR 0.225 ± 0.037 0.463 ± 0.057 0.143 ± 0.030 0.156 ± 0.031
Seq-MNAR 0.589 ± 0.026 0.218 ± 0.043 0.282 ± 0.041 0.285 ± 0.038
Polarization-MNAR 0.682 ± 0.033 0.318 ± 0.047 0.783 ± 0.027 0.741 ± 0.034
Soft-Polarization-MNAR 0.532 ± 0.040 0.391 ± 0.065 0.436 ± 0.033 0.331 ± 0.039
Latent-Factor-MNAR 0.468 ± 0.024 0.361 ± 0.042 0.106 ± 0.028 0.135 ± 0.027
Cluster-MNAR 0.452 ± 0.021 0.381 ± 0.042 0.059 ± 0.020 0.102 ± 0.024
Two-Phase-MNAR 0.555 ± 0.039 0.169 ± 0.043 0.409 ± 0.041 0.398 ± 0.039
Overall 0.482 ± 0.012 0.377 ± 0.017 0.252 ± 0.016 0.250 ± 0.015
Table 8: Mean Column-wise $R^{2}$ ± Standard Error by Missingness Pattern
Pattern EWF-TabPFN TabImpute HyperImpute ICE MissForest
MCAR 0.395 ± 0.036 0.414 ± 0.035 0.409 ± 0.035 0.387 ± 0.035 0.401 ± 0.036
NN-MNAR 0.387 ± 0.035 0.382 ± 0.034 0.364 ± 0.034 0.348 ± 0.033 0.351 ± 0.033
Self-Masking-MNAR 0.297 ± 0.039 0.285 ± 0.039 0.234 ± 0.038 0.250 ± 0.037 0.207 ± 0.030
Col-MAR 0.308 ± 0.033 0.289 ± 0.035 0.285 ± 0.034 0.287 ± 0.034 0.242 ± 0.032
Block-MNAR 0.283 ± 0.032 0.256 ± 0.033 0.242 ± 0.034 0.229 ± 0.033 0.221 ± 0.033
Seq-MNAR 0.234 ± 0.030 0.218 ± 0.027 0.246 ± 0.029 0.234 ± 0.030 0.210 ± 0.028
Panel-MNAR 0.275 ± 0.028 0.290 ± 0.029 0.285 ± 0.027 0.271 ± 0.029 0.271 ± 0.028
Polarization-MNAR 0.230 ± 0.022 0.190 ± 0.019 0.189 ± 0.020 0.177 ± 0.021 0.156 ± 0.017
Soft-Polarization-MNAR 0.177 ± 0.027 0.148 ± 0.024 0.146 ± 0.024 0.137 ± 0.023 0.131 ± 0.022
Latent-Factor-MNAR 0.347 ± 0.034 0.338 ± 0.033 0.335 ± 0.034 0.331 ± 0.033 0.322 ± 0.034
Cluster-MNAR 0.350 ± 0.036 0.342 ± 0.035 0.342 ± 0.037 0.325 ± 0.036 0.325 ± 0.036
Two-Phase-MNAR 0.370 ± 0.035 0.331 ± 0.037 0.333 ± 0.037 0.334 ± 0.036 0.310 ± 0.034
Censoring-MNAR 0.177 ± 0.025 0.143 ± 0.020 0.121 ± 0.019 0.125 ± 0.019 0.097 ± 0.017
Overall 0.295 ± 0.009 0.279 ± 0.009 0.272 ± 0.009 0.264 ± 0.009 0.249 ± 0.009
Pattern OT K-Nearest Neighbors MICE GAIN SoftImpute
MCAR 0.401 ± 0.034 0.323 ± 0.036 0.302 ± 0.036 0.318 ± 0.031 0.265 ± 0.041
NN-MNAR 0.334 ± 0.033 0.278 ± 0.031 0.254 ± 0.033 0.265 ± 0.026 0.223 ± 0.043
Self-Masking-MNAR 0.211 ± 0.034 0.207 ± 0.033 0.264 ± 0.044 0.224 ± 0.029 0.166 ± 0.030
Col-MAR 0.221 ± 0.031 0.227 ± 0.032 0.182 ± 0.030 0.156 ± 0.027 0.167 ± 0.033
Block-MNAR 0.199 ± 0.029 0.186 ± 0.029 0.155 ± 0.027 0.098 ± 0.016 0.165 ± 0.035
Seq-MNAR 0.192 ± 0.023 0.213 ± 0.027 0.139 ± 0.023 0.099 ± 0.015 0.181 ± 0.032
Panel-MNAR 0.247 ± 0.025 0.277 ± 0.028 0.146 ± 0.021 0.167 ± 0.023 0.181 ± 0.032
Polarization-MNAR 0.177 ± 0.018 0.146 ± 0.018 0.109 ± 0.015 0.109 ± 0.013 0.118 ± 0.021
Soft-Polarization-MNAR 0.108 ± 0.020 0.075 ± 0.018 0.066 ± 0.016 0.078 ± 0.014 0.077 ± 0.021
Latent-Factor-MNAR 0.306 ± 0.031 0.246 ± 0.031 0.227 ± 0.032 0.202 ± 0.023 0.194 ± 0.039
Cluster-MNAR 0.314 ± 0.033 0.254 ± 0.032 0.225 ± 0.032 0.239 ± 0.029 0.187 ± 0.038
Two-Phase-MNAR 0.260 ± 0.030 0.293 ± 0.031 0.191 ± 0.030 0.198 ± 0.026 0.190 ± 0.035
Censoring-MNAR 0.084 ± 0.016 0.095 ± 0.011 0.131 ± 0.022 0.127 ± 0.020 0.081 ± 0.021
Overall 0.235 ± 0.009 0.217 ± 0.008 0.184 ± 0.008 0.175 ± 0.007 0.169 ± 0.009
Pattern ForestDiffusion TabPFN MIWAE Col Mean
MCAR 0.220 ± 0.029 0.070 ± 0.008 0.036 ± 0.005 0.000 ± 0.000
NN-MNAR 0.177 ± 0.026 0.042 ± 0.006 0.025 ± 0.003 0.000 ± 0.000
Self-Masking-MNAR 0.116 ± 0.024 0.106 ± 0.016 0.026 ± 0.004 0.000 ± 0.000
Col-MAR 0.108 ± 0.020 0.081 ± 0.011 0.030 ± 0.003 0.000 ± 0.000
Block-MNAR 0.106 ± 0.017 0.051 ± 0.008 0.023 ± 0.003 0.000 ± 0.000
Seq-MNAR 0.089 ± 0.014 0.034 ± 0.003 0.022 ± 0.003 0.000 ± 0.000
Panel-MNAR 0.120 ± 0.016 0.083 ± 0.011 0.039 ± 0.006 0.000 ± 0.000
Polarization-MNAR 0.091 ± 0.012 0.036 ± 0.004 0.029 ± 0.004 0.000 ± 0.000
Soft-Polarization-MNAR 0.031 ± 0.004 0.004 ± 0.001 0.014 ± 0.001 0.000 ± 0.000
Latent-Factor-MNAR 0.152 ± 0.022 0.053 ± 0.007 0.028 ± 0.003 0.000 ± 0.000
Cluster-MNAR 0.181 ± 0.027 0.058 ± 0.009 0.033 ± 0.005 0.000 ± 0.000
Two-Phase-MNAR 0.126 ± 0.022 0.086 ± 0.012 0.022 ± 0.003 0.000 ± 0.000
Censoring-MNAR 0.071 ± 0.011 0.072 ± 0.018 0.067 ± 0.018 0.000 ± 0.000
Overall 0.122 ± 0.006 0.060 ± 0.003 0.030 ± 0.002 0.000 ± 0.000
Table 9: Mean Normalized Negative RMSE ± Standard Error by Missingness Pattern. Note that these normalized values are slightly different than Tab. 1 because the nonlinear model is included here.
[…]
MCAR 0.880 ± 0.021 0.803 ± 0.031 0.835 ± 0.022 0.866 ± 0.022
NN-MNAR 0.898 ± 0.019 0.758 ± 0.036 0.835 ± 0.022 0.816 ± 0.029
Self-Masking-MNAR 0.705 ± 0.037 0.689 ± 0.039 0.613 ± 0.043 0.639 ± 0.041
Col-MAR 0.868 ± 0.030 0.806 ± 0.040 0.836 ± 0.029 0.764 ± 0.038
Block-MNAR 0.890 ± 0.026 0.873 ± 0.027 0.868 ± 0.029 0.860 ± 0.023
Seq-MNAR 0.919 ± 0.015 0.863 ± 0.030 0.890 ± 0.018 0.828 ± 0.031
Panel-MNAR 0.893 ± 0.027 0.847 ± 0.034 0.448 ± 0.063 0.891 ± 0.022
Polarization-MNAR 0.800 ± 0.024 0.619 ± 0.039 0.751 ± 0.030 0.564 ± 0.032
Soft-Polarization-MNAR 0.756 ± 0.045 0.709 ± 0.034 0.808 ± 0.034 0.665 ± 0.041
Latent-Factor-MNAR 0.888 ± 0.020 0.773 ± 0.035 0.857 ± 0.020 0.832 ± 0.024
Cluster-MNAR 0.896 ± 0.016 0.833 ± 0.029 0.858 ± 0.017 0.824 ± 0.022
Two-Phase-MNAR 0.879 ± 0.026 0.859 ± 0.031 0.871 ± 0.026 0.867 ± 0.020
Censoring-MNAR 0.711 ± 0.032 0.790 ± 0.039 0.623 ± 0.039 0.673 ± 0.036
Overall 0.845 ± 0.008 0.786 ± 0.010 0.776 ± 0.010 0.776 ± 0.009
@@
Dataset TabImpute EWF-TabPFN HyperImpute MissForest
EgyptianSkulls 0.901 0.917 0.961 0.946
FacultySalaries 0.607 0.693 0.811 0.567
Glass-Classification 0.889 0.937 0.810 0.886
ICU 1.059 1.045 1.137 1.112
MercuryinBass 1.311 1.373 1.377 1.343
Rainfall-in-Kerala-1901-2017 0.908 0.944 0.902 0.950
SMSA 0.840 0.814 0.939 0.786
SolarPower 0.889 0.805 0.873 0.888
Student-Scores 0.429 0.392 0.441 0.435
Swiss-banknote-conterfeit-detection 0.774 0.976 0.856 0.730
WineDataset 0.829 0.928 0.803 0.797
alcohol-qcm-sensor 0.467 0.590 0.580 0.517
analcatdata_election2000 0.364 0.538 0.662 0.615
analcatdata_gviolence 0.687 0.796 0.801 0.877
analcatdata_olympic2000 1.950 1.871 1.922 1.895
appendicitis_test 0.694 0.861 0.650 0.792
appendicitis_test_edsa 0.534 0.668 0.532 0.569
autoPrice 0.735 0.834 0.762 0.702
baskball 1.008 1.028 1.100 1.109
benzo32 1.070 1.055 1.055 1.053
breast-cancer-coimbra 1.068 1.058 1.183 1.098
chatfield_4 0.473 0.476 0.531 0.492
chscase_vine1 0.928 0.809 0.996 0.833
dataset_analcatdata_creditscore 1.238 1.063 1.131 1.122
divorce_prediction 0.542 0.525 0.528 0.467
edm 0.612 0.763 0.553 0.566
humans_numeric 0.989 0.952 0.970 0.991
machine_cpu 0.902 1.039 0.956 0.858
metafeatures 2.609 2.263 2.090 2.225
pollution 1.946 1.967 2.065 1.837
pwLinear 1.031 1.075 1.356 1.356
pyrim 0.847 0.877 0.863 0.859
rabe_131 1.016 0.792 1.003 1.031
rabe_148 0.905 0.985 1.017 0.958
rabe_265 1.104 1.131 1.298 1.090
sleuth_case1201 0.940 1.054 0.939 0.963
sleuth_ex1605 1.026 1.087 1.257 1.071
slump 0.750 0.871 0.650 0.805
treepipit 1.226 1.265 1.169 1.172
visualizing_hamster 0.754 0.872 0.905 0.750
wisconsin 0.854 0.777 0.685 0.783
witmer_census_1980 0.654 0.704 0.616 0.637
Table 13: Non-normalized RMSE Values for MCAR Pattern by Dataset including ReMasker. Note that we could not replicate ReMasker’s results using their open-source implementation. In order to not misrepresent their method, we do not include this table in our main results.
Dataset TabImpute HyperImpute MissForest ReMasker
EgyptianSkulls 0.901 0.961 0.946 2.329
FacultySalaries 0.607 0.811 0.567 1.534
Glass-Classification 0.889 0.810 0.886 3.126
ICU 1.059 1.137 1.112 1.479
MercuryinBass 1.311 1.377 1.343 2.198
Rainfall-in-Kerala-1901-2017 0.908 0.902 0.950 3.560
SMSA 0.840 0.939 0.786 1.979
SolarPower 0.889 0.873 0.888 1.589
Student-Scores 0.429 0.441 0.435 1.785
Swiss-banknote-conterfeit-detection 0.774 0.856 0.730 1.364
WineDataset 0.829 0.803 0.797 1.454
alcohol-qcm-sensor 0.467 0.580 0.517 2.044
analcatdata_election2000 0.364 0.662 0.615 4.563
analcatdata_gviolence 0.687 0.801 0.877 1.617
analcatdata_olympic2000 1.950 1.922 1.895 2.787
appendicitis_test 0.694 0.650 0.792 2.174
appendicitis_test_edsa 0.534 0.532 0.569 1.986
autoPrice 0.735 0.762 0.702 2.190
baskball 1.008 1.100 1.109 2.246
benzo32 1.070 1.055 1.053 3.282
breast-cancer-coimbra 1.068 1.183 1.098 1.820
chatfield_4 0.473 0.531 0.492 3.545
chscase_vine1 0.928 0.996 0.833 2.720
dataset_analcatdata_creditscore 1.238 1.131 1.122 1.469
divorce_prediction 0.542 0.528 0.467 1.216
edm 0.612 0.553 0.566 1.415
humans_numeric 0.989 0.970 0.991 4.165
machine_cpu 0.902 0.956 0.858 1.402
metafeatures 2.609 2.090 2.225 2.715
pollution 1.946 2.065 1.837 2.985
pwLinear 1.031 1.356 1.356 1.040
pyrim 0.847 0.863 0.859 1.647
rabe_131 1.016 1.003 1.031 2.076
rabe_148 0.905 1.017 0.958 2.508
rabe_265 1.104 1.298 1.090 2.633
sleuth_case1201 0.940 0.939 0.963 2.455
sleuth_ex1605 1.026 1.257 1.071 2.327
slump 0.750 0.650 0.805 1.960
treepipit 1.226 1.169 1.172 1.648
visualizing_hamster 0.754 0.905 0.750 1.884
wisconsin 0.854 0.685 0.783 2.422
witmer_census_1980 0.654 0.616 0.637 2.487